OpenStudy (anonymous):

a particular hybrid car travels approximately 208 miles on 4 gal of gas. find the amount of gas required for 936 mi trip?

OpenStudy (anonymous):

936 / 208 = 4.5....4.5time more than 208 so then 4.5 * 4 is your answer.

OpenStudy (anonymous):

use the ratio of 4/208 and g/936 4/208=g/936 936(1/52)=g g=18